# Precipitation Downscaling with Spatiotemporal Video Diffusion

### v2

@@ Abstract
In climate science and meteorology, high-resolution local precipitation (rain and snowfall) predictions are limited by the computational costs of simulation-based methods. Statistical downscaling, or super-resolution, is a common workaround where a low-resolution prediction is improved using statistical approaches. Unlike traditional computer vision tasks, weather and climate applications require capturing the accurate conditional distribution of high-resolution given low-resolution patterns to assure reliable ensemble averages and unbiased estimates of extreme events, such as heavy rain. This work extends recent video diffusion models to precipitation super-resolution, employing a deterministic downscaler followed by a temporally-conditioned diffusion model to capture noise characteristics and high-frequency patterns. We test our approach on FV3GFS output, an established large-scale global atmosphere model, and compare it against five state-of-the-art baselines. Our analysis, capturing CRPS, MSE, precipitation distributions, and qualitative aspects using California and the Himalayas as examples, establishes our method as a new standard for data-driven precipitation downscaling.
@@ 1 Introduction
Patterns of precipitation (rainfall and snowfall) are central to human and natural life on earth. In a rapidly warming climate, reliable simulations of how precipitation patterns may change can help human societies adapt to climate change. However, such simulations are challenging as weather systems vary over many scales in space and time. Moreover, interactions with complex surface features such as mountains and coastlines contribute to precipitation trends and extremes [48]. For many purposes, such as estimating flood hazards, precipitation must be estimated at a spatial resolutions of only a few kilometers. Fluid-dynamical models of the global atmosphere are prohibitively expensive to run at such fine scales [55], so the climate adaptation community relies on “downscaling” [^1] This is the climate science terminology for super-resolution. of coarse-grid simulations to a finer grid. Traditional downscaling methods are either “dynamical” (running a fine-grid fluid-dynamical model limited to the region of interest, which requires specialized knowledge and computational resources) or “statistical” (typically restricted to simple univariate methods) [57]. Advances in super-resolution methods developed by the computer vision community are poised to dramatically improve statistical downscaling.
Such work is a natural follow-up to recent advancements in deep-learning-based weather and climate prediction methodologies, which have ushered in a new era of data-driven forecasting. These approaches show remarkable performance gains, boasting improvements of orders of magnitude in runtime without sacrificing accuracy [37, 24].
@@
Our focus centers on addressing a temporal downscaling problem. Our objective is to enhance the resolution of a sequence (“video”) of low-resolution precipitation frames into a sequence of high-resolution frames. Precipitation has strong temporal continuity on hourly time scales. Hence, despite the apparent differences between precipitation data and natural videos, we adopt ideas from video super-resolution to leverage information from multiple context frames to stochastically downscale the precipitation sequence [43, 33].
@@
Figure 1: Illustration of our model’s training and inference pipelines. Blue blocks are common to both phases, red blocks are exclusive to training, and green blocks are for inference only. The model takes a low-resolution precipitation sequence, downscales it deterministically using spatio-temporal factorized attention, and models residual with conditional diffusion and factorized attention. Here, $T$ denotes the sequence length, and $N$ denotes the diffusion steps. The model parameters ($\theta={\phi,\psi}$) are jointly optimized during training. See Appendix A.1 for further details on our model architecture.
@@
Until recently, endeavors to enhance the resolution of climate states like precipitation have been primarily based on deterministic regression methods that employ convolutions or transformers. However, super-resolution is a one-to-many mapping with a continuum of “correct” answers. Treating such one-to-many problems using supervised learning is therefore prone to visual artifacts since the supervised network tries to predict the average from a multitude of incompatible solutions. Such mode averaging is a known source of blurriness in visual data [25, 66]. Besides visual artifacts, mode averaging can have even more dramatic implications in climate and weather modeling, such as the underestimation of extreme precipitation, which is mainly induced by regional weather patterns on the unresolved scale. While some successful supervised super-resolution methods for images and video have been proposed [9, 22, 65, 8, 20], a natural approach to preventing mode averaging are conditional generative models, due to their ability to capture multi-modal conditional data distributions.
@@
To that end, a line of recent work proposes the use of generative adversarial networks (GANs) for precipitation downscaling. These GAN-based methodologies often encounter challenges, displaying a tendency to converge on a specific mode of the data distribution, occasionally fixating on isolated points in extreme cases. Despite their perceptual appeal, the scientific utility of super-resolution necessitates the accurate modeling of the statistical distribution of high-resolution data given low-resolution input, a facet that GANs typically fall short of capturing.
@@
In this paper, we propose Spatiotemporal Video Diffusion (STVD), a diffusion-based generative model for spatio-temporal precipitation downscaling. Our model uses a spatio-temporal factorized attention-based downscaler to incorporate temporal context, followed by modeling the residual error with a similar factorized attention-based conditional diffusion model, allowing us to add finer details to the coarse prediction made by the downscaler. Diffusion models are particularly well-suited for precipitation downscaling, as they have been shown to successfully capture high dimensional and multimodal distributions, alleviating a key drawback of GAN-based methods for climate science applications.
@@
This study underscores the capability of conditional diffusion models to meet the specific needs of statistical precipitation downscaling. Specifically, our contributions are:
@@
    We introduce a novel framework for temporal precipitation downscaling using diffusion models. Our model combines a deterministic downscaling module with a diffusion-based residual module, leveraging spatiotemporal, factorized attention to process information from multiple low-resolution frames.
@@
    Our model outperforms five strong super-resolution baselines across multiple criteria, including MSE and several distributional metrics. We compare against two image super-resolution models and three video super-resolution models using the FV3GFS global atmosphere simulation dataset [67, 7].
@@
Our paper is structured as follows. First, we describe our method (Sec. 2), which is then followed by a discussion on our experimental findings (Sec. 3). Lastly, we discuss relevant literature (Sec. 4) and its connection to our work.
@@ 2.0.1 Problem Statement
At training time, we assume that we have access to a collection of high-resolution precipitation frame sequences $\mathbf{y}^{0:T}$ and their corresponding low-resolution precipitation frame sequences $\mathbf{x}^{0:T}$. Such a low-resolution sequence can be obtained through area-weighted coarsening [34] of the corresponding high-resolution sequence. The dataset is discussed extensively in 3.1. Frame indices are represented by superscripts, where we assume that each sequence consists of $T+1$ frames for simplicity. We note that it may be possible to roll-out predictions for multiple sequences autoregressively using techniques such as reconstruction guidance [18], but we leave exploration of this direction for future work. Our objective is to train a model to effectively downscale, or super-resolve a given sequence $\mathbf{x}^{0:T}$ with $\mathbf{y}^{0:T}$ serving as the target. Note we use both “downscaling” and “super-resolution” interchangeably.
More formally, let $\mathbf{x}^{t}\in\mathbb{R}^{C\times H\times W}$ and $\mathbf{y}^{t}\in\mathbb{R}^{1\times sH\times sW}$ represent individual low-resolution and high-resolution frames. Here, $s\in\mathbb{N}$ denotes the downscaling factor, $C$ represents the number of climate states used as input to the model, and $H,W$ indicate the height and width of the low-resolution frame, respectively. For our study, we adopt an downscaling factor of $s=8$ and have $C=12$ total low-resolution channels. In addition to the low-resolution precipitation state, we provide an additional eleven channels of information to the model via concatenation, such as topography, wind velocity, and surface temperature; see Appendix A.2 for details.
Figure 2: A qualitative comparison between our proposed model and baselines for a precipitation event associated with a cold front impinging on the Northern California coast and then the Sierra mountain range (coastline marked in hazy white). Fig. 4 plots the regional topography. The time interval between adjacent frames is 3 hours; the plotted region is $1000\times 1000$ km. Our model resolves the fine-grid precipitation structure better than the considered baselines. See Appendix A.3 for additional samples.
@@ 2.0.2 Solution Sketch
We approach the downscaling problem as a conditional generative modeling task. Accordingly, we devise a model with the objective of learning conditional distribution of high-resolution precipitation frames, taking into account the contextual information provided by the low-resolution precipitation frame sequence.
Our proposed solution, Spatiotemporal Video Diffusion (STVD) (Fig. 1), relies on two modules: a deterministic downscaler based on spatio-temporal factorised attention and a stochastic component based on conditional diffusion models [17, 53], also leveraging factorised attention. The first module uses a UNet with spatio-temporal attention to integrate information from a low-resolution frame sequence. This process results in a high-resolution frame sequence, denoted by $\bar{\mathbf{y}}^{0:T}$, which serves as an initial prediction and is further refined by the second module. The second module is a conditional diffusion model that stochastically generates a sequence of additive residual frames $\mathbf{r}^{0:T}$ which serves to add fine-grained details to the initial prediction. In all, composing these two modules results in a predicted high-resolution frame sequence of the form $\hat{\mathbf{y}}^{0:T}=\bar{\mathbf{y}}^{0:T}+\mathbf{r}^{0:T}$. Both modules are trained end-to-end.
Decomposing the prediction target into a deterministic mean and a stochastic residual is inspired by recent findings in predictive-coding-based video decompression. Here, one aims to predict (in the best way possible) a sequence of video frames while compressing the sparse residuals [2, 64] that can be modeled more easily than dense frames. Similarly, it is easier to generate residuals than dense images when using diffusion models [63, 35].
In what follows, we first describe our overall probabilistic framework for downscaling. Then, we discuss the deterministic module along with temporal attention, followed by the remaining residual prediction module based on diffusion generative modeling. See Appendix A.1 for the details of our model architecture.
@@ 2.1 Probabilistic Modeling of Downscaling
Given a sequence of low-resolution frames $\mathbf{x}^{0:T}$ and the corresponding sequence of high-resolution frames $\mathbf{y}^{0:T}$, we seek to learn a parametric approximation $p_{\theta}$ of the underlying conditional distribution $p\left(\mathbf{y}^{0:T}\mid\mathbf{x}^{0:T}\right)\approx p_{\theta}\left(\mathbf{y}^{0:T}\mid\mathbf{x}^{0:T}\right)$. Importantly, we make no independence assumptions across time, i.e. each generated frame $\mathbf{y}^{t}$ is allowed to depend on all other generated frames. Moreover, the generated high-resolution frame sequence is conditioned on the entire low-resolution frame sequence. This allows us to capture any long-range temporal correlations between the frames, thereby enhancing the fidelity and cohesion of the high-resolution frame reconstruction.
As discussed in the overview of this section, the likelihood $p_{\theta}\left(\mathbf{y}^{0:T}\mid\mathbf{x}^{0:T}\right)$ is modeled using a combination of a deterministic downscaler and a residual diffusion model. We will discuss how the model parameters $\theta=(\phi,\psi)$ decompose into those for a downscaler ($\phi$) and a diffusion model ($\psi$).
@@ 2.1.1 Deterministic Downscaling
Our first module is a deterministic downscaler that predicts an initial high-resolution frame sequence
$\bar{\mathbf{y}}^{0:T}=\mu_{\phi}\left({\mathbf{x}}^{0:T}\right),$ (1)
where $\mu_{\phi}$ is a network generating a deterministic high-resolution prediction with parameters $\phi$. We choose to perform bicubic interpolation on each frame of $\mathbf{x}^{0:T}$ before passing the frame sequence through the network $\mu_{\phi}$. As the diffusion network necessarily operates on high-resolution inputs (i.e. denoising the high-resolution residuals), this choice allows us to use the same UNet [42] architecture (with different weights) for both the downscaling module $\mu_{\phi}$ and the residual diffusion module. This enables us to easily share features across the modules via concatenation. See Appendix A.1 for further details.
Importantly, $\mu_{\phi}$ incorporates a temporal attention mechanism that allows any given frame at time $t$, or its corresponding feature map, to attend to all the context frames from 00 to $T$. This mechanism’s architecture allows for the concurrent inference of all frames within the sequence $\bar{\mathbf{y}}^{0:T}$. We emphasize that these attention weights are different for each frame, allowing for the flexible incorporation of information across time.
@@ 2.1.2 Stochastic Residual Modeling via Diffusion
After computing an initial high-resolution sequence $\bar{\mathbf{y}}^{0:T}$, the finer details that refine this prediction are modeled by residuals learned via a conditional diffusion model. That is, our final stochastic high-resolution frame sequence $\hat{\mathbf{y}}^{0:T}$ is generated by sampling an additive residual sequence $\mathbf{r}^{0:T}$ from a conditional diffusion model, i.e.
$\hat{\mathbf{y}}^{0:T}=\bar{\mathbf{y}}^{0:T}+\mathbf{r}^{0:T}.$ (2)
Thus, we seek to model the residuals $\mathbf{r}^{0:T}=\mathbf{y}^{0:T}-\mathbf{\bar{y}}^{0:T}$. We emphasize that our diffusion model generates the entire residual sequence $\mathbf{r}^{0:T}$ concurrently, with the generative process of residual $\mathbf{r}^{t}$ dependent on the other residuals in the sequence. This is achieved via a UNet architecture with spatio-temporal attention, similar to the mechanism used for the deterministic downscaling module.
In order to model the distribution of $\mathbf{r}^{0:T}$, we leverage the framework of diffusion models, in particular the DDPM model [17]. To that end, we introduce a collection of latent variables $\mathbf{r}_{0:N}^{0:T}$, where the lower subscripts indicate the denoising diffusion step. In the forward process of a diffusion model, the latent variable $\mathbf{r}_{n}^{0:T}$ is created from $\mathbf{r}_{n-1}^{0:T}$ via additive noise. In the reverse process, which serves to perform generation, a denoising model (with parameters $\psi$) is trained to predict $\mathbf{r}_{j-1}^{0:T}$ from $\mathbf{r}_{j}^{0:T}$.
We use $N$ to denote the total number of denoising steps. Note that $\mathbf{r}^{0:T}=\mathbf{r}_{0}^{0:T}$, i.e. the first diffusion step, corresponds to the true residual. We further note that $\mathbf{r}_{0}^{0:T}$ implicitly depends on the downscaler parameters $\phi$, and hence we may simultaneously optimize all model parameters $\theta=(\phi,\psi)$ within the context of diffusion modeling.
As is standard in diffusion models [17], we parameterize the reverse process via a Gaussian distribution whose mean is determined by a neural network $M_{\psi}$,
@@
where $M_{\psi}$ is a denoising network (a UNet with spatio-temporal attention). Here, the variance $\gamma$ is a hyperparameter. We note that the diffusion model directly accesses the context $\mathbf{c}=({\mathbf{x}}^{0:T},\overline{\mathbf{y}}^{0:T})$, but is also implicitly conditioned on ${\mathbf{x}}^{0:T}$ via concatenation of feature maps from the downscaler module. As in the downscaler, we bicubically upsample $\mathbf{x}^{0:T}$ before channel-wise concatenation with $\overline{\mathbf{y}}^{0:T}$ to match the dimensions when forming $\mathbf{c}$.
while not converged do        Sample $\mathbf{x}^{0:T}$ and $\mathbf{y}^{0:T}$;        $n\sim\mathcal{U}(0,1,2,..,N)$;        $\epsilon\sim\mathcal{N}(\textbf{0},\textbf{I})$;        $\bar{\mathbf{y}}^{0:T}=\mu_{\phi}\left({\mathbf{x}}^{0:T}\right)$;        $\mathbf{r}^{0:T}_{0}=\mathbf{y}^{0:T}-\bar{\mathbf{y}}^{0:T}$;        $\mathbf{v}=\alpha_{n}\epsilon-\sigma_{n}\mathbf{r}^{0:T}_{0}$;        $\mathbf{r}_{n}^{0:T}=\alpha_{n}\mathbf{r}_{0}^{0:T}+\sigma_{n}\epsilon$;        $\mathbf{c}=\left(\mathbf{x}^{0:T},\overline{y}^{0:T}\right)$;        $\hat{\mathbf{v}}=M_{\psi}\left(\mathbf{r}_{n}^{0:T},n,\mathbf{c}\right)$;        $L=||\mathbf{v}-\hat{\mathbf{v}}||^{2}$;        $(\psi,\phi)=(\psi,\phi)-\nabla_{\psi,\phi}L$; Algorithm 1 Training STVD Get an equally spaced increasing sub-sequence $\tau$ of $\left[1,\ldots,N\right]$ of length $S\ll N$, i.e., $\{\tau_{1},\ldots,\tau_{S}\}$; $\bar{\mathbf{y}}^{0:T}=\mu_{\phi}\left(\mathbf{x}^{0:T}\right)$; $\mathbf{c}=\left(\mathbf{x}^{0:T},\overline{y}^{0:T}\right)$; $\mathbf{r}_{S}^{0:T}\sim\mathcal{N}(\textbf{0},\textbf{I})$; for n in reversed($\tau$) do        $\hat{\mathbf{v}}=M_{\psi}\left(\mathbf{r}_{n}^{0:T},n,\mathbf{c}\right)$;        $\hat{\mathbf{r}}=\alpha_{n}\mathbf{r}_{n}^{0:T}-\sigma_{n}\hat{\mathbf{v}}$;        $\hat{\epsilon}=\frac{\sigma_{n}}{\alpha_{n}}\left(\mathbf{r}_{n}^{0:T}-\hat{\mathbf{r}}\right)$;        $\mathbf{r}^{0:T}_{n-1}=\alpha_{n-1}\hat{\mathbf{r}}+\sigma_{n-1}\hat{\epsilon}$; $\hat{\mathbf{y}}^{0:T}=\bar{\mathbf{y}}^{0:T}+\mathbf{r}^{0:T}_{0}$; Algorithm 2 Sampling STVD

### v3

@@ Abstract
In climate science and meteorology, high-resolution local precipitation (rain and snowfall) predictions are limited by the computational costs of simulation-based methods. Statistical downscaling, or super-resolution, is a common workaround where a low-resolution prediction is improved using statistical approaches. Unlike traditional computer vision tasks, weather and climate applications require capturing the accurate conditional distribution of high-resolution given low-resolution patterns to assure reliable ensemble averages and unbiased estimates of extreme events, such as heavy rain. This work extends recent video diffusion models to precipitation super-resolution, employing a deterministic downscaler followed by a temporally-conditioned diffusion model to capture noise characteristics and high-frequency patterns. We test our approach on FV3GFS output, an established large-scale global atmosphere model, and compare it against six state-of-the-art baselines. Our analysis, capturing CRPS, MSE, precipitation distributions, and qualitative aspects using California and the Himalayas as examples, establishes our method as a new standard for data-driven precipitation downscaling.
@@ 1 Introduction
Precipitation patterns are central to human and natural life. In a rapidly warming climate, reliable simulations of changing precipitation patterns can help adapt to climate change. However, these simulations are challenging due to the multi-scale variability of weather systems and the influence of complex surface features (like mountains and coastlines) on precipitation trends and extremes [53]. For many purposes, such as estimating flood hazards, precipitation must be estimated at spatial resolutions of only a few kilometers. Fluid-dynamical models of the global atmosphere are too expensive to run routinely at such fine scales [60], so the climate adaptation community relies on “downscaling” [^1] This is the climate science terminology for super-resolution. of coarse-grid simulations to a finer grid. Traditional downscaling methods are either “dynamical” (running a fine-grid fluid-dynamical model limited to the region of interest, which requires specialized knowledge and computational resources) or “statistical” (typically restricted to simple univariate methods) [62]. Our work builds on vision based super-resolution methods to improve statistical downscaling and is a natural follow-up to recent deep-learning-based weather/climate prediction methods, which have revolutionized data-driven forecasting. These approaches boast improvements of orders of magnitude in runtime without sacrificing accuracy [41, 27].
@@
We address the downscaling problem for a sequence. Our objective is to transform a sequence (“video”) of low-resolution precipitation frames into a sequence of high-resolution frames. Despite differences from natural videos, precipitation’s hourly temporal continuity allows us to use video super-resolution techniques to leverage multiple context frames for stochastic downscaling [48, 34].
@@
Figure 1: Our model’s training and inference pipelines: Blue blocks apply to both phases, red blocks to training only, and green blocks to inference only. It deterministically downscales a low-resolution precipitation sequence using spatio-temporal factorized attention and models residuals with conditional diffusion (with factorized attention). Here, $T$ denotes sequence length and $N$ denotes diffusion steps. The parameters ($\theta={\phi,\psi}$) are optimized jointly during training. See A.1 for details.
@@
Recent efforts to enhance the resolution of climate states like precipitation have relied on deterministic regression methods using convolutions or transformers. However, super-resolution is a one-to-many mapping with a continuum of “correct” answers. Supervised learning for these problems often leads to visual artifacts from mode averaging, where the network predicts an average of incompatible solutions, causing blurriness in visual data [28, 71]. Besides visual artifacts, mode averaging can have even more dramatic implications in climate and weather modeling, such as the underestimation of extreme precipitation, which is mainly induced by regional weather patterns on the unresolved scale. A natural alternative to supervised super-resolution methods [11, 25, 70, 10, 21] to prevent mode averaging is conditional generative modeling, which captures multimodal conditional distributions.
@@
To that end, recent works propose using generative adversarial networks (GANs) for precipitation downscaling. These methods often face challenges, tending to converge on specific modes of the data distribution and occasionally fixating on isolated points in extreme cases. Despite their perceptual appeal, the scientific utility of super-resolution requires accurate modeling of the statistical distribution of high-resolution data given low-resolution input, which GANs typically fail to capture.
@@
In this paper, we propose SpatioTemporal Video Diffusion (STVD), for spatio-temporal precipitation downscaling. It uses a downscaler for a coarse prediction. The residual error is modeled with conditional diffusion, adding finer details to the coarse prediction. Both modules rely on spatio-temporal factorized attention to process the input sequence. Diffusion models are well-suited for precipitation downscaling as they successfully capture high dimensional and multimodal distributions, alleviating a key drawback of GAN-based methods for climate science applications.
@@
This study highlights the capability of conditional diffusion models to meet the specific needs of statistical precipitation downscaling, with our key contributions being:
@@
    We introduce a novel framework for temporal precipitation downscaling using diffusion models. Our model combines a deterministic downscaling module with a diffusion-based residual module. It leverages spatio-temporal factorized attention to process information from multiple low-resolution frames.
@@
    Our model outperforms six strong super-resolution baselines across multiple criteria, including MSE and several distributional metrics. We compare against two image super-resolution models and four video super-resolution models using the FV3GFS global atmosphere simulation dataset [72, 9].
@@
Our paper is structured as follows: we first describe our method (Sec. 2), followed by our experimental findings (Sec. 3). Finally, we discuss relevant literature (Sec. 4) and its connection to our work.
@@ 2.0.1 Problem Statement
At training time, we assume access to a collection of high-resolution precipitation frame sequences $\mathbf{y}^{0:T}$ and their corresponding low-resolution precipitation frame sequences $\mathbf{x}^{0:T}$. Such a low-resolution sequence can be obtained through area-weighted coarsening [35] of the corresponding high-resolution sequence. The dataset is discussed extensively in Sec. 3. Frame indices are represented by superscripts, where we assume that each sequence consists of $T+1$ frames for simplicity. While it may be possible to roll out predictions for multiple sequences autoregressively using techniques such as reconstruction guidance [19], we leave this exploration for future work. Our objective is to train a model to effectively downscale, or super-resolve a given sequence $\mathbf{x}^{0:T}$ with $\mathbf{y}^{0:T}$ serving as the target. We use “downscaling” and “super-resolution” interchangeably.
More formally, let $\mathbf{x}^{t}\in\mathbb{R}^{C\times H\times W}$ and $\mathbf{y}^{t}\in\mathbb{R}^{1\times sH\times sW}$ represent individual low-resolution and high-resolution frames. Here, $s\in\mathbb{N}$ denotes the downscaling factor, $C$ is the number of channels (quantities used as input to the model to characterize the atmospheric state in each low-resolution grid cell so as to add skill to the precipitation prediction), and $H,W$ indicate the height and width of the low-resolution frame. For our study, we adopt a downscaling factor of $s=8$ and have $C=12$ total low-resolution channels. In addition to the low-resolution precipitation state, we provide eleven channels of information to the model, such as topography, wind velocity, and surface temperature; see A.2 for details.
Figure 2: A qualitative comparison between our proposed model and top baseline for a precipitation event associated with a cold front impinging on the Northern California coast and then the Sierra mountain range (coastline marked in hazy white). Fig. 5 plots the regional topography. The time interval between adjacent frames is 3 hours; the plotted region is $1000\times 1000$ km. Our model resolves the fine-grid precipitation structure better than the considered baselines. See A.3 for full-page high quality samples from Himalayas and Sierra.
@@ 2.0.2 Solution Sketch
Our approach treats the downscaling problem as a conditional generative modeling task. We devise a model to learn the conditional distribution of high-resolution precipitation frames, incorporating contextual information from the low-resolution precipitation frame sequence.
Our proposed solution, SpatioTemporal Video Diffusion (STVD) (Fig. 1), relies on two modules: a deterministic downscaler and a stochastic component based on conditional diffusion models [18, 58], both using spatio-temporal factorized attention. The first module uses a UNet with factorized attention to integrate information from a low-resolution frame sequence, resulting in an initial prediction frame sequence. The second module is a conditional diffusion model that stochastically generates a sequence of additive residual frames $\mathbf{r}^{0:T}$ which serves to add fine-grained details to the initial prediction. Together, these two modules produce a high-resolution frame sequence $\hat{\mathbf{y}}^{0:T}=\bar{\mathbf{y}}^{0:T}+\mathbf{r}^{0:T}$. Both modules are trained end-to-end.
Decomposing the prediction into a deterministic mean and a stochastic residual is inspired by predictive-coding-based video decompression. This approach aims to predict a sequence of video frames while compressing the sparse residuals [1, 69] which are easier to model than dense frames. Similarly, it is easier to generate residuals than dense images when using diffusion models [68, 37].
In what follows, we first describe our overall probabilistic framework for downscaling. Then, we discuss the deterministic module along with spatio-temporal attention, followed by the remaining residual prediction module based on diffusion generative modeling. See A.1 for architecture details.
@@ 2.1 Probabilistic Modeling of Downscaling
Given a sequence of low-resolution frames $\mathbf{x}^{0:T}$ and the corresponding high-resolution frames $\mathbf{y}^{0:T}$, we aim to learn a parametric approximation $p_{\theta}$ of the conditional distribution $p\left(\mathbf{y}^{0:T}\mid\mathbf{x}^{0:T}\right)\approx p_{\theta}\left(\mathbf{y}^{0:T}\mid\mathbf{x}^{0:T}\right)$. Importantly, we do not assume independence across time; each generated frame $\mathbf{y}^{t}$ can depend on all other generated frames. The generated high-resolution frame sequence is conditioned on the entire low-resolution frame sequence, capturing long-range temporal correlations and enhancing the fidelity and cohesion of the high-resolution reconstruction.
As noted earlier, the likelihood $p_{\theta}\left(\mathbf{y}^{0:T}\mid\mathbf{x}^{0:T}\right)$ is modeled using a deterministic downscaler and a residual diffusion model. We will discuss how the model parameters $\theta=(\phi,\psi)$ decompose into those for a downscaler ($\phi$) and a diffusion model ($\psi$).
@@ 2.1.1 Deterministic Downscaling
Our first module is a deterministic downscaler that predicts an initial high-resolution frame sequence $\bar{\mathbf{y}}^{0:T}=\mu_{\phi}\left({\mathbf{x}}^{0:T}\right)$ where $\mu_{\phi}$ is a network generating a deterministic high-resolution prediction with parameters $\phi$. We perform bicubic interpolation on each frame of $\mathbf{x}^{0:T}$ before passing the sequence through the network $\mu_{\phi}$. Since the diffusion network operates on high-resolution inputs (i.e. denoising the high-resolution residuals), this choice allows us to use the same UNet [47] architecture (with different weights) for both the downscaling module $\mu_{\phi}$ and the residual diffusion module. This enables us to easily share features across the modules via concatenation. See A.1 for further details.
Importantly, $\mu_{\phi}$ incorporates a temporal attention mechanism that allows any frame at time $t$, or its corresponding feature map, to attend to all context frames from 00 to $T$. This architecture enables the concurrent inference of all frames within the sequence $\bar{\mathbf{y}}^{0:T}$. The attention weights differ for each frame, allowing for the flexible incorporation of information across time.
@@ 2.1.2 Stochastic Residual Modeling via Diffusion
After computing the initial prediction $\bar{\mathbf{y}}^{0:T}$, finer details are modeled by residuals learned from a conditional diffusion model. Our final stochastic high-resolution frame sequence $\hat{\mathbf{y}}^{0:T}$ is generated by sampling an additive residual sequence $\mathbf{r}^{0:T}$ from this model: $\hat{\mathbf{y}}^{0:T}=\bar{\mathbf{y}}^{0:T}+\mathbf{r}^{0:T}.$ Thus, we seek to model the residuals $\mathbf{r}^{0:T}=\mathbf{y}^{0:T}-\mathbf{\bar{y}}^{0:T}$. Our diffusion model generates the entire residual sequence $\mathbf{r}^{0:T}$ concurrently, with the generation of each residual $\mathbf{r}^{t}$ dependent on the others. This is achieved via a UNet architecture with spatio-temporal attention, similar to the mechanism used for the deterministic downscaling module. See A.1 for further details.
To model the distribution of $\mathbf{r}^{0:T}$, we use DDPM [18]. To that end, we introduce a collection of latent variables $\mathbf{r}_{0:N}^{0:T}$, where the lower subscripts indicate the denoising diffusion step. In the forward process, the latent variable $\mathbf{r}_{n}^{0:T}$ is created from $\mathbf{r}_{n-1}^{0:T}$ via additive noise. In the reverse process for generation, a denoising model (with parameters $\psi$) is trained to predict $\mathbf{r}_{j-1}^{0:T}$ from $\mathbf{r}_{j}^{0:T}$. $N$ denotes the total number of denoising steps. Note that $\mathbf{r}^{0:T}=\mathbf{r}_{0}^{0:T}$, i.e. the first diffusion step corresponds to the true residual. Additionally, $\mathbf{r}_{0}^{0:T}$ implicitly depends on the downscaler parameters $\phi$, allowing us to simultaneously optimize all model parameters $\theta=(\phi,\psi)$ within the context of diffusion modeling.
As is standard in diffusion models [18], we parameterize the reverse process via a Gaussian distribution with a mean determined by a neural network $M_{\psi}$,
@@
where $M_{\psi}$ is a denoising network and $\gamma$ is a hyperparameter for variance. The diffusion model directly accesses the context $\mathbf{c}=({\mathbf{x}}^{0:T},\overline{\mathbf{y}}^{0:T})$, and is implicitly conditioned on ${\mathbf{x}}^{0:T}$ via concatenation of feature maps from the downscaler module. As in the downscaler, we bicubically upsample $\mathbf{x}^{0:T}$ before channel-wise concatenation with $\overline{\mathbf{y}}^{0:T}$ to match the dimensions when forming $\mathbf{c}$.
[Uncaptioned image] Figure 3: Tradeoff between mean square error and percentile error (see Sec. 3). Inference at Himalayan region (see Figs. 5 and 10). while not converged do        Sample $\mathbf{x}^{0:T}$ and $\mathbf{y}^{0:T}$;        $n\sim\mathcal{U}(0,1,2,..,N)$;        $\epsilon\sim\mathcal{N}(\textbf{0},\textbf{I})$;        $\bar{\mathbf{y}}^{0:T}=\mu_{\phi}\left({\mathbf{x}}^{0:T}\right)$;        $\mathbf{r}^{0:T}_{0}=\mathbf{y}^{0:T}-\bar{\mathbf{y}}^{0:T}$;        $\mathbf{v}=\alpha_{n}\epsilon-\sigma_{n}\mathbf{r}^{0:T}_{0}$;        $\mathbf{r}_{n}^{0:T}=\alpha_{n}\mathbf{r}_{0}^{0:T}+\sigma_{n}\epsilon$;        $\mathbf{c}=\left(\mathbf{x}^{0:T},\overline{y}^{0:T}\right)$;        $\hat{\mathbf{v}}=M_{\psi}\left(\mathbf{r}_{n}^{0:T},n,\mathbf{c}\right)$;        $L=||\mathbf{v}-\hat{\mathbf{v}}||^{2}$;        $(\psi,\phi)=(\psi,\phi)-\nabla_{\psi,\phi}L$; Algorithm 1 Training STVD Get an equally spaced increasing sub-sequence $\tau$ of length $K\ll N$; $\bar{\mathbf{y}}^{0:T}=\mu_{\phi}\left(\mathbf{x}^{0:T}\right)$; $\mathbf{c}=\left(\mathbf{x}^{0:T},\overline{y}^{0:T}\right)$; $\mathbf{r}_{K}^{0:T}\sim\mathcal{N}(\textbf{0},\textbf{I})$; for n in reversed($\tau$) do        $\hat{\mathbf{v}}=M_{\psi}\left(\mathbf{r}_{n}^{0:T},n,\mathbf{c}\right)$;        $\hat{\mathbf{r}}=\alpha_{n}\mathbf{r}_{n}^{0:T}-\sigma_{n}\hat{\mathbf{v}}$;        $\hat{\epsilon}=\frac{\sigma_{n}}{\alpha_{n}}\left(\mathbf{r}_{n}^{0:T}-\hat{\mathbf{r}}\right)$;        $\mathbf{r}^{0:T}_{n-1}=\alpha_{n-1}\hat{\mathbf{r}}+\sigma_{n-1}\hat{\epsilon}$; $\hat{\mathbf{y}}^{0:T}=\bar{\mathbf{y}}^{0:T}+\mathbf{r}^{0:T}_{0}$; Algorithm 2 Sampling STVD